OpenStudy (anonymous):

1. find the equation of the tangent line to f(x)= 3x+1, x=4 which I got the first part to find P (4,13) 2. find the equation of the tangent line f(x)= sqrt (x) , x= 1/9 which I got the first part P ( 1/9, 1/3) 3. find the equation to the tangent line f(x)= 1/sqrt(x), x=1 .Again got P(1, 1)

OpenStudy (campbell_st):

can you find the 1st derivative in question 1..?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

and for 2. I have 1 / 2 sqrt(x) and for 3 I have 1/(2*x^(3/2))

OpenStudy (campbell_st):

ok... so for 2 substitute x = 1/9 will give the slope at x = 1/9 then find the equation of the straight line .. for 3 substitute x = 1 and this will be the slope of the tangent... then you can use the point to find the equation of the line...

OpenStudy (campbell_st):

for the 1st question the slope is 3 and the point is (4, 13) you'll see something interesting when you find the equation of the tangent.

OpenStudy (anonymous):

for question 2 I got y= 3/2x + 3/18

OpenStudy (campbell_st):

using y = 3/2x + b 1/3 = 3/2(1/9) + b so 1/3 = 1/6 + b so b = 1/6 ...

OpenStudy (anonymous):

question 1 is 3x+1 which is the same from the original problem

OpenStudy (campbell_st):

and actually there is another tangent in question 2, where m = -3/2 so y = -3/2 x + b 1/3 = -3/2(1/9) + b 1/2 = b

OpenStudy (campbell_st):

thats correct...

OpenStudy (campbell_st):

but thats interesting since a tangent only touches at 1 point... the equation you have touches at all points..

OpenStudy (campbell_st):

in question 2 \[\sqrt{\frac{1}{9}} = \pm \frac{1}{3} \]

OpenStudy (campbell_st):

I think in question 3 your 1st derivative is incorrect \[y = \frac{1}{\sqrt{x} }.... or......y = x^{-\frac{1}{2}}\]

OpenStudy (campbell_st):

so \[\frac{dy}{dx} = - \frac{1}{2} x^{-\frac{3}{2}} ... or ... \frac{dy}{dx} = -\frac{ 1}{2\sqrt{x^3}}\]

OpenStudy (campbell_st):

so again 2 tangents... as you have 2 slopes...

OpenStudy (anonymous):

I get y= -1/2x + 3/2 for the equation

OpenStudy (campbell_st):

that looks good... I have to go... hope I helped...

OpenStudy (anonymous):

thank you